2. An object is thrown straight upwards at a speed of 25 m/sec. How high will it rise? How fast will it be moving when s it is 20 m high? How long will it take to reach this height on + the way up? On the way down?
Answer #1

2. at highest point, its speed will be zero.

vi = 25 m/s , vf =0 and a = - 9.8m/s^2

using vf^2 - vi^2 = 2 a d

0^2 - 25^2 = 2(-9.8)(H)

H = 31.9 m ..........Ans


v^2 - 25^2 = 2(-9.8)(20)

v = 15.3 m/s ........Ans

While going up,

vf = vi + a t

15.3 = 25 - 9.8 t

t = 1 sec .......Ans

while going down, vf =- 15.3 m/s

- 15.3 = 25 - 9.8 t

t = 4.11 sec ......Ans